OpenStudy (anonymous):

i will medal if its right please help me @horsegirl27 @HelpBlahBlahBlah @highjumpcollect @mickeyboo07

OpenStudy (anonymous):

here

OpenStudy (kainui):

Any ideas?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do u know it

OpenStudy (kainui):

Remember, x is just a number, but we don't know what it is yet, so simple things like this: \[\LARGE 3*(3)=9\] Have to get changed, since we can rewrite 3=5-2 this is the same kind of idea, we're just distributing \[\LARGE 3(5-2)=3*5-3*2=15-6=9\]

OpenStudy (anonymous):

so is it A, B, C, or D

OpenStudy (anonymous):

@Kainui @King.Void.

OpenStudy (anonymous):

@Nnesha please help

OpenStudy (king.void.):

I'm not good at Math. Sorry!